# Invariance transformations in wave-optics lensing: implications for gravitational-wave astrophysics and cosmology

### v1

@@ IV.1 Mismatch
Thanks to the normalization, if two signals are identical, the match is 1. On the contrary, the mismatch is defined as $\epsilon=1-$match. The mismatch is an interesting statistic for two main reasons. It measures the differences between two signals independently of their amplitude, or SNR $\rho=\sqrt{(h|h)}$. Moreover, in the limit of a high-SNR signal in some data, the mismatch $\epsilon$ can be related to the improvement in the $\chi^{2}$ goodness-of-fit to the data of one model versus the other, $\Delta\chi^{2}\approx 2\epsilon\rho^{2}$ (see e.g. Ezquiaga et al. (2022)). Therefore, a high mismatch indicates that it would be possible to clearly prefer one hypothesis over the other.
Because lensed waveforms have an arrival time delay, they need to be aligned in time first to maximize the match between them. To align the lensed waveforms for different lens parameters, the amplification factor needs to be multiplied by a time shift term $\exp(-iwt_{\lambda})$, which is expanded as
@@
The term $\exp(-iw\lambda^{2}y^{2}/2)$ in the time alignment cancels the term $\exp(iw\lambda^{2}y^{2}/2)$ in the analytical solution of $F_{\lambda}$ in Eq. (33), so that only the terms with $w\lambda$ are left in $F_{\lambda}$, which makes it impossible to distinguish $M_{L}$ and $\lambda$ through comparing GW waveforms since $w=8\pi(1+z_{L})M_{L}f$. Therefore if one were to align the waveforms with parameter-dependent time delay in Eq. (44) with exact values of $M_{L}$ and $\lambda$, the MSD could not be broken for the PM model. We illustrate the degeneracy in the upper panel of Figure 1, where the mismatch is computed using the LIGO A+ design sensitivity curve Barsotti et al. (2018). There is a band of zero mismatch that tracks the expected $M_{L}=M_{L,\mathrm{fid}}/\lambda$ scaling over the whole range of $\lambda$, where the fiducial values are $M_{L,\mathrm{fid}}=700M_{\odot}$ and $\lambda_{\mathrm{fid}}=1$.
However, in the actual observations, the exact time delay of an image cannot be obtained by using Eq. (33), since the lens parameters $M_{L}$, $y$ or $\lambda$ are not known. The only observable from the GW signals is the coalescence time $t_{c}$. Therefore, when performing parameter estimation (which we will show in section IV.3), the waveform cannot be shifted by a different time delay for each parameter-space point. Instead, we compute the mismatch between the signals shifted by the same amount of time for all parameter-space points. We choose this fiducial time shift to be the value of the time delay of the fiducial waveform. Then we shift all waveforms constructed with various $M_{L}$ and $\lambda$ with this fiducial time shift. As shown in the lower panel of Figure 1, the MSD is now broken in the mismatch plot. The region with low mismatch still follows $M_{L}\propto 1/\lambda$ but it is now bounded. This indicates that the MSD can be broken in parameter estimation for lensed waveforms. Moreover, we learn that the correlation between the arrival time and the waveform distortion is key in breaking this degeneracy.
Figure 1: Mismatch between time-aligned lensed waveforms for different $M_{L}$ and $\lambda$ and the injected lensed waveform for $M_{L}=700\,M_{\odot}$ and $\lambda=1$. The injected parameters for the source are listed in Table 1 (left column). The green dashed line indicates $M_{L}=M_{L,\mathrm{fid}}/\lambda$. Upper panel: The match is aligned with time shifts computed with Eq. (26) for each pair of $M_{L}$ and $\lambda$. Lower panel: The match is aligned with the time shift for the fiducial $M_{L}$ and $\lambda$.
@@ IV.2 Fisher Matrix
For the next analysis, we assume that the noise is stationary and Gaussian and that we are in the limit of high SNR, so that the likelihood peak becomes narrow enough that the prior around the maximum-likelihood value $\theta_{ML}^{i}$ can be taken as approximately flat. Then the posterior distribution of the parameters $\theta^{i}=\theta_{ML}^{i}+\Delta\theta^{i}$ given the data $d$ is obtained by Finn (1992); Cutler and Flanagan (1994)
@@
We plot the contours for the Fisher matrix constraints on the parameter set $\{{\cal M}_{c},q,d_{L},t_{c},M_{L},y,\lambda\}$ in Figure 2, where ${\cal M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass; $q=m_{2}/m_{1}$ is the mass ratio of the binary source with component masses $m_{1}$ and $m_{2}$ ($m_{2}<m_{1}$); $d_{L}$ is the luminosity distance of the source; $M_{L}$ is the mass of the lens; and $y$ is the impact parameter. Here, and in the whole paper, we refer to the detector-frame parameters, so, for example, $M_{L}$ stands for the mass of the lens with redshift applied from the detector frame. The injected values for the binary source parameters are listed in Table 1 (left column), which is the same as in Figure 1. We used the LIGO A+ design sensitivity curve Barsotti et al. (2018) to compute the Fisher matrix. As shown in Figure 2, the contours of the Fisher matrix forecast for several pairs of parameters are very elongated, indicating strong degeneracy between each pair of them. The ellipses are however finite, indicating as well that the degeneracy is broken at some point. The MSD is shown by the contour for $(M_{L},\lambda)$, which matches the results shown by the mismatch analysis in Figure 1. Besides, the luminosity distance $d_{L}$ and the coalescence time $t_{c}$ are both degenerate with $M_{L}$ and $\lambda$. This is because the amplification factor is proportional to $1/\lambda$ as shown in Eq. (33), and the waveform is proportional to $1/d_{L}$, and thus $d_{L}$ is degenerate with $\lambda$. And because of the degeneracy between $M_{L}$ and $\lambda$, $d_{L}$ is also degenerate with $M_{L}$. On the other hand, $t_{c}$ shifts the waveform in time by multiplying the waveform by $\exp(2\pi ft_{c})$. However, the term $\exp(i\omega\lambda^{2}y^{2}/2)$ in the amplification factor in Eq. (33) gives an equivalent effect to a time shift, so it creates degeneracy between $t_{c}$ and $M_{L}$, and $t_{c}$ and $\lambda$.
@@ IV.3 Bayesian Parameter Estimation
We present three main cases: (i) an injected event with high SNR and a PM lens model; (ii) an injected event with high SNR and an SIS lens model; and (iii) an event similar to the real GW200208_130117 Abbott et al. (2023b) (from here on GW200208 for short). For the first two cases, we injected a regular and a biased signal, i.e. with $\lambda=1$ and $\lambda\neq 1$, respectively. The details of the parameters used in these cases can be found in Table 1, and more information about the software used can be found in Appendix B.
@@ IV.3.3 GW200208-like event
While both previous cases have high SNRs, we present now a more typical scenario for current ground-based detectors. This analysis is done on an injection that follows the characteristics of the real event GW200208333N.B.: the real event is not considered to be lensed. We use here the parameters that give the most similar waveform to GW200208 if it were lensed., that are listed in Tab. 1. As the lens mass is still rather small, we decide to use the PM lens model both for injection and PE analysis. This event gives us the possibility to see how the PE and MSD behave in a scenario that is more realistic given the current detected events.
@@ VI.2 Wave-optics cosmography
We first generate GW events following the Powerlaw + Peak black-hole mass distribution model Fishbach et al. (2020); Farah et al. (2022); Abbott et al. (2023c), as was also done in the LVK cosmology paper Abbott et al. (2023d). We adopt the mass model parameters as $\alpha=3.78$, $\beta=0.81$, $m_{\rm min}^{\rm BH}=4.98M_{\odot}$, $m_{\rm max}^{\rm BH}=112.5M_{\odot}$, $\delta_{m}=4.8M_{\odot}$, $\mu_{g}=32.27M_{\odot}$, $\sigma_{g}=3.88M_{\odot}$ and $\lambda_{p}=0.03$, based on detected LVK events Abbott et al. (2023c). We use the GWCOSMO package Gray et al. (2020, 2022, 2023) to generate the Powerlaw + Peak mass distribution, which is shown in Figure 23 in Appendix C.
Following Ref. Callister et al. (2020), we use the Madau-Dickinson star formation rate redshift evolution function Madau and Dickinson (2014) to describe the redshift distribution of the binary merger rate, which gives
@@
We adopt the same value $R_{0}=30\leavevmode\nobreak\ {\rm yr}^{-1}{\rm Gpc}^{-1}$ as in Callister et al. (2020), which is consistent with the GWTC-3 population study Abbott et al. (2023c). We also adopt $\gamma=4.59$, $k=2.86$ and $z_{p}=2.47$ for the Madau-Dickinson model as in Abbott et al. (2023d).
Next, we calculate the rate of microlensing for GW events in terms of the optical depth. We assume that the lens mass density is constant in comoving volume and takes up $10\%$ of the matter density of the universe. The differential optical depth is given in Vietri and Ostriker (1983); Zumalacarregui and Seljak (2018) as
@@
where $\alpha=0.1$ is the fraction of matter density contributing to lens density, and we adopt the 2018 Planck TT+lensing+BAO value $\Omega_{m,0}=0.3065$ Aghanim et al. (2018). By integrating Eq. (49) over $z_{L}$ up to the source redshift, we obtain the optical depth of lensing effects for GW sources, which is plotted in Figure 24 in Appendix C. The lensed GW event rate as a function of redshift is then calculated by multiplying the merger rate redshift evolution function with the optical depth in terms of redshift. A plot showing the lensed event rate compared with the binary merger rate is presented in Figure 25 in Appendix C. Note that the fraction of dark matter made of these lenses does not affect the redshift dependence of the optical depth and only rescales the overall rate. This fraction should be chosen not to conflict with current constraints on massive compact halo objects (e.g. Mróz et al. (2024)). In our case, we are interested mostly in simulating lensed events and not in their relative rate compared to non-lensed events.
For the next step, we draw lensed GW events with random masses and redshifts with the probability weighted by the mass distribution and the lensing rate probability distribution computed above. The redshift range is $z<1$ since LVK sources at high redshift are likely to have SNR lower than the detection threshold (unless strongly lensed). We select events with SNR $>8$. Then we make very simple assumptions that each lens weights $100\,M_{\odot}$, and the lenses are distributed uniformly in comoving volume, which gives
@@
where $E(z)=\sqrt{\Omega_{m,0}(1+z)^{3}+\Omega_{\Lambda,0}}$, and $\Omega_{\Lambda,0}=1-\Omega_{m,0}$. So we draw the redshift $z_{L}$ of the corresponding lens for each event randomly from a uniform-in-comoving-volume distribution below the source redshift. In addition, we assign a random value for the impact parameter $y$ from a uniform distribution in $(0,10)$ for each lens. Then we construct the lensed waveform under the transformation with $h_{0}$, and forecast the constraints on $h_{0}$. We simulate the uncertainties in $h_{0}$ via FM for 100 lensed events with the LIGO A+ design sensitivity curve Barsotti et al. (2018) for one LIGO detector, and show them in Figure 10 along with the redshifted source chirp masses and SNRs.
Figure 10: The plot shows 100 simulated lensed GW events in the wave-optics regime detected by a ground-based GW detector with LIGO A+ design sensitivity Barsotti et al. (2018). The lens mass is assumed to be $100\,M_{\odot}$ for each event. The horizontal axis shows the redshifted chirp mass of the simulated events. The vertical axis shows the uncertainty in $h_{0}$ for each event. The colour scale shows the SNR of each event.
Figure 10 shows that the uncertainties in $h_{0}$ from Fisher matrix forecasts for wave-optics lensed waveforms for PM lenses of $100\,M_{\odot}$ are mostly on the order of unity or larger. This means that the uncertainty of constraints on $H_{0}$ is mostly larger than 100%, which is equivalent to unconstrained results. Such uncertainties are extremely large compared to other $H_{0}$ measurements, which are currently reaching the percent level precision (see the review in Abdalla et al. (2022)). With LVK detections, we are not expected to have a large enough number of wave-optics lensed events to achieve a precise $H_{0}$ measurement. However, next-generation (3G) ground-based detectors, such as the Einstein Telescope (ET) Hild et al. (2011); ET Steering Committee Editorial Team (2020); Maggiore et al. (2020) or Cosmic Explorer Reitze et al. (2019); Evans et al. (2021), are expected to detect significantly more lensed GW events per year Li et al. (2018). So we also forecast $h_{0}$ uncertainties with the ET sensitivity curve Hild et al. (2011) in Figure 11.
Figure 11: The plot shows 100 simulated lensed GW events in the wave-optics regime similar to Figure 10, but detected by the Einstein Telescope with sensitivity as in Hild et al. (2011).
@@
In summary, this forecast presents a possibility to measure $H_{0}$ with wave-optics lensed GW events by 3G detectors. However, our forecast constraints are based on our assumption of a fixed lens mass. We will leave a more accurate forecast with a more realistic lens population model and lensed GW detection rate for 3G detectors for future work.

### v2

@@ IV.1 Mismatch
Thanks to the normalization, if two signals are identical, the match is 1. On the contrary, the mismatch is defined as $\epsilon=1-$match. The mismatch is an interesting statistic for two main reasons. It measures the differences between two signals independently of their amplitude, or SNR $\rho=\sqrt{(h|h)}$. Moreover, in the limit of a high-SNR signal in some data, the mismatch $\epsilon$ can be related to the improvement in the $\chi^{2}$ goodness-of-fit to the data of one model versus the other, $\Delta\chi^{2}\approx 2\epsilon\rho^{2}$ (see e.g. [88]). Therefore, a high mismatch indicates that it would be possible to clearly prefer one hypothesis over the other.
Because lensed waveforms have an arrival time delay, they need to be aligned in time first to maximize the match between them. For the lensed waveforms we generate at different points in the ($M_{L}$, $\lambda$) space while fixing other parameters, the only time differences between each waveform is this arrival time delay due to lensing, so the maximization of matches over time shift can be achieved by removing this time delay computed at each ($M_{L}$, $\lambda$) point. To do so, the amplification factor needs to be multiplied by a time shift term $\exp(-iwt_{\lambda})$, which is expanded using Eq. (26) in dimensionless unit as
@@
The term $\exp(-iw\lambda^{2}y^{2}/2)$ in the time alignment cancels the term $\exp(iw\lambda^{2}y^{2}/2)$ in the analytical solution of $F_{\lambda}$ in Eq. (33), so that only the terms with $w\lambda$ are left in $F_{\lambda}$, which makes it impossible to distinguish $M_{L}$ and $\lambda$ through comparing GW waveforms since $w=8\pi(1+z_{L})M_{L}f$. Therefore if one were to align the waveforms with parameter-dependent time delay in Eq. (44) with exact values of $M_{L}$ and $\lambda$, the MSD could not be broken for the PM model. We illustrate the degeneracy in the upper panel of Figure 1, where the mismatch is computed using the LIGO A+ design sensitivity curve [89]. There is a band of zero mismatch that tracks the expected $M_{L}=M_{L,\mathrm{fid}}/\lambda$ scaling over the whole range of $\lambda$, where the fiducial values are $M_{L,\mathrm{fid}}=700M_{\odot}$ and $\lambda_{\mathrm{fid}}=1$.
However, in the actual observations, the exact time delay of an image cannot be obtained by using Eq. (33), since the lens parameters $M_{L}$, $y$ or $\lambda$ are not known. The only observable from the GW signals is the coalescence time $t_{c}$. Therefore, when performing parameter estimation (which we will show in section IV.3), the waveform cannot be shifted by a time delay computed by values of ($M_{L}$, $\lambda$) at different parameter-space points. Instead, we compute the mismatch between the signals shifted by the same amount of time for all parameter-space points to illustrate the scenario where the exact time shift is unknown. We choose this fiducial time shift to be the value of the time delay of the fiducial waveform. Then we shift all waveforms constructed with various ($M_{L}$, $\lambda$) with this fiducial time shift. As shown in the lower panel of Figure 1, the MSD is now broken in the mismatch plot. The region with low mismatch still follows $M_{L}\propto 1/\lambda$ but it is now bounded. Although theoretically one can search for a value of $t_{c}$ that maximizes the matches over time shift, it now also depends on parameters other than ($M_{L}$, $\lambda$). This shows that the correlation between the arrival time and the waveform distortion is key in breaking this degeneracy. As we will see in Section IV.3, in parameter estimation for detected signals in a multi-parameter space, the degeneracy appears between $t_{c}$ and lensing parameters, but the degeneracy between $M_{L}$ and $\lambda$ becomes bounded.
Figure 1: Mismatch between time-aligned lensed waveforms for different $M_{L}$ and $\lambda$ and the injected lensed waveform for $M_{L}=700\,M_{\odot}$ and $\lambda=1$. The injected parameters for the source are listed in Table 1 (left column). The green dashed line indicates $M_{L}=M_{L,\mathrm{fid}}/\lambda$. Upper panel: The match is aligned with time shifts computed with Eq. (26) at each ($M_{L}$, $\lambda$) point. Lower panel: The match is aligned with a fixed time shift computed with the fiducial $M_{L}$ and $\lambda$.
@@ IV.2 Fisher Matrix
For the next analysis, we assume that the noise is stationary and Gaussian and that we are in the limit of high SNR, so that the likelihood peak becomes narrow enough that the prior around the maximum-likelihood value $\theta_{ML}^{i}$ can be taken as approximately flat. Then the posterior distribution of the parameters $\theta^{i}=\theta_{ML}^{i}+\Delta\theta^{i}$ given the data $d$ is obtained by [90, 91]
@@
We plot the contours for the Fisher matrix constraints on the parameter set $\{{\cal M}_{c},q,d_{L},t_{c},M_{L},y,\lambda\}$ in Figure 2, where ${\cal M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass; $q=m_{2}/m_{1}$ is the mass ratio of the binary source with component masses $m_{1}$ and $m_{2}$ ($m_{2}<m_{1}$); $d_{L}$ is the luminosity distance of the source; $M_{L}$ is the mass of the lens; and $y$ is the impact parameter. Here, and in the whole paper, we refer to the detector-frame parameters, so, for example, $M_{L}$ stands for the mass of the lens with redshift applied from the detector frame. The injected values for the binary source parameters are listed in Table 1 (left column), which is the same as in Figure 1. We used the LIGO A+ design sensitivity curve [89] to compute the Fisher matrix. As shown in Figure 2, the contours of the Fisher matrix forecast for several pairs of parameters are very elongated, indicating strong degeneracy between each pair of them. The ellipses are however finite, indicating as well that the degeneracy is broken at some point. The MSD is shown by the contour for $(M_{L},\lambda)$, which matches the results shown by the mismatch analysis in Figure 1. Besides, the luminosity distance $d_{L}$ and the coalescence time $t_{c}$ are both degenerate with $M_{L}$ and $\lambda$. This is because the amplification factor is proportional to $1/\lambda$ as shown in Eq. (33), and the waveform is proportional to $1/d_{L}$, and thus $d_{L}$ is degenerate with $\lambda$. And because of the degeneracy between $M_{L}$ and $\lambda$, $d_{L}$ is also degenerate with $M_{L}$. On the other hand, $t_{c}$ shifts the waveform in time by multiplying the waveform by $\exp(2\pi ft_{c})$. However, the term $\exp(i\omega\lambda^{2}y^{2}/2)$ in the amplification factor in Eq. (33) gives an equivalent effect to a time shift, so it creates degeneracy between $t_{c}$ and $M_{L}$, and $t_{c}$ and $\lambda$.
@@ IV.3 Bayesian Parameter Estimation
We present three main cases: (i) an injected event with high SNR and a PM lens model; (ii) an injected event with high SNR and an SIS lens model; and (iii) an event similar to the real GW200208_130117 [92] (from here on GW200208 for short). For the first two cases, we injected a regular and a biased signal, i.e. with $\lambda=1$ and $\lambda\neq 1$, respectively. The details of the parameters used in these cases can be found in Table 1, and more information about the software used can be found in Appendix B.
@@ IV.3.3 GW200208-like event
While both previous cases have high SNRs, we present now a more typical scenario for current ground-based detectors. This analysis is done on an injection that follows the characteristics of the real event GW200208333N.B.: The real event is not considered to be lensed, but it raised interest in the community as a possible lensed candidate. We use here the parameters that give the most similar waveform to GW200208 if it were lensed., that are listed in Tab. 1. As the lens mass is still rather small, we decide to use the PM lens model both for injection and PE analysis. This event gives us the possibility to see how the PE and MSD behave in a scenario that is more realistic given the current detected events.
@@ VI.2 Wave-optics cosmography
We first generate GW events following the Powerlaw + Peak black-hole mass distribution model [94, 95, 96], as was also done in the LVK cosmology paper [97]. We adopt the mass model parameters as $\alpha=3.78$, $\beta=0.81$, $m_{\rm min}^{\rm BH}=4.98M_{\odot}$, $m_{\rm max}^{\rm BH}=112.5M_{\odot}$, $\delta_{m}=4.8M_{\odot}$, $\mu_{g}=32.27M_{\odot}$, $\sigma_{g}=3.88M_{\odot}$ and $\lambda_{p}=0.03$, based on detected LVK events [96]. We use the GWCOSMO package [98, 99, 100] to generate the Powerlaw + Peak mass distribution, which is shown in Figure 23 in Appendix C.
Following Ref. [101], we use the Madau-Dickinson star formation rate redshift evolution function [102] to describe the redshift distribution of the binary merger rate, which gives
@@
We adopt the same value $R_{0}=30\leavevmode\nobreak\ {\rm yr}^{-1}{\rm Gpc}^{-1}$ as in [101], which is consistent with the GWTC-3 population study [96]. We also adopt $\gamma=4.59$, $k=2.86$ and $z_{p}=2.47$ for the Madau-Dickinson model as in [97].
Next, we calculate the rate of microlensing for GW events in terms of the optical depth. We assume that the lens mass density is constant in comoving volume and takes up $10\%$ of the matter density of the universe (This is an arbitrary choice, as this work only provides a proof of principle). The differential optical depth is given in [103, 104] as
@@
where $\alpha=0.1$ is the fraction of matter density contributing to lens density, and we adopt the 2018 Planck TT+lensing+BAO value $\Omega_{m,0}=0.3065$ [105]. By integrating Eq. (49) over $z_{L}$ up to the source redshift, we obtain the optical depth of lensing effects for GW sources, which is plotted in Figure 24 in Appendix C. The lensed GW event rate as a function of redshift is then calculated by multiplying the merger rate redshift evolution function with the optical depth in terms of redshift. A plot showing the lensed event rate compared with the binary merger rate is presented in Figure 25 in Appendix C. Note that the fraction of dark matter made of these lenses does not affect the redshift dependence of the optical depth and only rescales the overall rate. This fraction should be chosen not to conflict with current constraints on massive compact halo objects (e.g. [106]). In our case, we are interested mostly in simulating lensed events and not in their relative rate compared to non-lensed events.
For the next step, we draw lensed GW events with random masses and redshifts with the probability weighted by the mass distribution and the lensing rate probability distribution computed above. The redshift range is $z<1$ since LVK sources at high redshift are likely to have SNR lower than the detection threshold (unless strongly lensed). We select events with SNR $>8$. Then we make a simple assumption that each lens weights $100\,M_{\odot}$, because we want to ensure that the lensed GW events are in the wave-optics regime, since LVK BBH event total masses are around 10 to 100 solar masses. And we want to focus on analysing the factors from the sources that affect constraints on $h_{0}$ by fixing lens masses. We will also investigate the effect of changing lens masses later. In addition, we assume that the lenses are distributed uniformly in comoving volume, which gives
@@
where $E(z)=\sqrt{\Omega_{m,0}(1+z)^{3}+\Omega_{\Lambda,0}}$, and $\Omega_{\Lambda,0}=1-\Omega_{m,0}$. So we draw the redshift $z_{L}$ of the corresponding lens for each event randomly from a uniform-in-comoving-volume distribution below the source redshift. In addition, we assign a random value for the impact parameter $y$ from a uniform distribution in $(0,10)$ for each lens. Then we construct the lensed waveform under the transformation with $h_{0}$ with the PM lens model, and forecast the constraints on $h_{0}$. Given the lens mass allowed to cause wave-optics lensing for LVK events, the lenses are likely stellar-origin or intermediate black holes, so that they can be well described by the PM lens model. We simulate the uncertainties in $h_{0}$ via FM for 100 lensed events with the LIGO A+ design sensitivity curve [89] for one LIGO detector, and show them in Figure 10 along with the redshifted source chirp masses and SNRs.
Figure 10: The plot shows 100 simulated lensed GW events in the wave-optics regime detected by a ground-based GW detector with LIGO A+ design sensitivity [89]. The lens mass is assumed to be $100\,M_{\odot}$ for each event. The horizontal axis shows the redshifted chirp mass of the simulated events. The vertical axis shows the uncertainty in $h_{0}$ for each event. The colour scale shows the SNR of each event.
Figure 10 shows that the uncertainties in $h_{0}$ from Fisher matrix forecasts for wave-optics lensed waveforms for PM lenses of $100\,M_{\odot}$ are mostly on the order of unity or larger. This means that the uncertainty of constraints on $H_{0}$ is mostly larger than 100%, which is equivalent to unconstrained results. Such uncertainties are extremely large compared to other $H_{0}$ measurements, which are currently reaching the percent level precision (see the review in [107]). With LVK detections, we are not expected to have a large enough number of wave-optics lensed events to achieve a precise $H_{0}$ measurement. However, next-generation (3G) ground-based detectors, such as the Einstein Telescope (ET) [108, 109, 110] or Cosmic Explorer [111, 112], are expected to detect significantly more lensed GW events per year [31]. So we also forecast $h_{0}$ uncertainties with the ET sensitivity curve [108] in Figure 11.
Figure 11: The plot shows 100 simulated lensed GW events in the wave-optics regime similar to Figure 10, but detected by the Einstein Telescope with sensitivity as in [108].
@@
In summary, this forecast presents a possibility to measure $H_{0}$ with wave-optics lensed GW events by 3G detectors. By combining constraints on $h_{0}$ from multiple lensed events, the uncertainty can drop to a much lower level. However, readers should be aware that our forecast constraints are based on our assumption of a simple lens population model and a fixed lens mass, so our forecasts may not represent the realistic precision for future detections. The current work mainly stands as a proof of principle. We will leave a more accurate forecast with a more realistic lens population model and lensed GW detection rate for 3G detectors to future works.